OpenStudy (anonymous):

The sum of the first "n" terms of a series is Sn = n^2 + 4n. Find t1, t2, t3

OpenStudy (shamim):

ok

OpenStudy (shamim):

u wrote\[s _{n}=n ^{2}+4n\]

OpenStudy (shamim):

sorry

OpenStudy (shamim):

need time

OpenStudy (anonymous):

kk np

OpenStudy (raden):

t1 = a1 = S1 put n=1 to Sn

OpenStudy (shamim):

i think the series will b 5+7+9+.....................+(2n+3)

OpenStudy (shamim):

so t1=5 t2=7 t3=9

OpenStudy (anonymous):

Can I get explanations on how to do it? o.O Not just the answers?

OpenStudy (shamim):

if u put n=1 u will get 2n+3=2*1+3=5

OpenStudy (shamim):

so t1=5

OpenStudy (anonymous):

Why do you put n=1? How do you know to make it 1 and not any other number? o.O

OpenStudy (shamim):

when n=1 we will get the first term of the series

OpenStudy (shamim):

when we will put n=2 in (2n+3) then we will get the second term of the series

OpenStudy (anonymous):

wait if u put in the second term doesnt that equal the first term plus the second term?

OpenStudy (shamim):

no

OpenStudy (shamim):

when u will put n=2 in the equation \[s _{n}=n ^{2}+4n\]

OpenStudy (anonymous):

because the equation Sn = n^2 + 4n is for the sum and n is the terms in the series. So if you put 1 in you'll get the first term because it doesnt get added to anything, but if you put 2 terms in the equation you will get t1 + t2

OpenStudy (shamim):

then u will get the sum of 1st term and second term

OpenStudy (anonymous):

Yea okay lol thats wat i meant

OpenStudy (anonymous):

so youd have to subtract the previous terms to get the one ur looking for. Correct?

OpenStudy (shamim):

ya

OpenStudy (anonymous):

Kay! Thank you! :)

OpenStudy (shamim):

thanks for medal

OpenStudy (shamim):

welcome

OpenStudy (anonymous):

Np you deserved it ^_^

OpenStudy (shamim):

anyway may i know the name of ur country and class

OpenStudy (anonymous):

United States and I'm taking Pre-calculus right now

OpenStudy (shamim):

which class u r reading

OpenStudy (anonymous):

?

OpenStudy (shamim):

i mean r u a student of Advance level or something else

OpenStudy (anonymous):

Um... I don't know what you call it from where you're from, but I'm taking Honors Pre-Calc which is like the more advanced course of Pre Calc here ^_^

OpenStudy (shamim):

understood

OpenStudy (shamim):

i m frm bangladesh

OpenStudy (anonymous):

I see! Nice to meet you ^_^

OpenStudy (shamim):

nicve to meet u

OpenStudy (shamim):

nice to meet u